# Exploring Behavior-Relevant and Disentangled Neural Dynamics with Generative Diffusion Models

### v1

@@ Abstract
Understanding the neural basis of behavior is a fundamental goal in neuroscience. Current research in large-scale neuro-behavioral data analysis often relies on decoding models, which quantify behavioral information in neural data but lack details on behavior encoding. This raises an intriguing scientific question: “how can we enable in-depth exploration of neural representations in behavioral tasks, revealing interpretable neural dynamics associated with behaviors”. However, addressing this issue is challenging due to the varied behavioral encoding across different brain regions and mixed selectivity at the population level. To tackle this limitation, our approach, named “BeNeDiff”, first identifies a fine-grained and disentangled neural subspace using a behavior-informed latent variable model. It then employs state-of-the-art generative diffusion models to synthesize behavior videos that interpret the neural dynamics of each latent factor. We validate the method on multi-session datasets containing widefield calcium imaging recordings across the dorsal cortex. Through guiding the diffusion model to activate individual latent factors, we verify that the neural dynamics of latent factors in the disentangled neural subspace provide interpretable quantifications of the behaviors of interest. At the same time, the neural subspace in BeNeDiff demonstrates high disentanglement and neural reconstruction quality. Our codes are available at https://github.com/yulewang97/BeNeDiff.
@@ 1 Introduction
Understanding and elucidating the complex interrelationships between behavioral data and neural population activity is a long-standing goal in systems neuroscience (Batty et al., 2019; Gomez-Marin et al., 2014; Krakauer et al., 2017; Berman, 2018). Exploring the neural basis of behavior not only deepens our basic knowledge of brain functions but also establishes a foundation for developing improved treatments for psychiatric and neurological conditions (Vieira et al., 2017; Ibáñez et al., 2018). Significant progress has been achieved in developing computational toolkits for neuro-behavioral decoding by using behavior video data (Whiteway et al., 2021; Batty et al., 2019; Musall et al., 2019; Stringer et al., 2019). These methods perform region-based behavior decoding to map regions of neural activity among the entire dorsal cortex to the behaviors from the videos. However, these methods only quantify how much behavioral information is encoded in neural populations, but do not reveal the details of such encoding. There has been markedly less focus, with cortex-wide signals, on enabling in-depth exploration of neural activities during behavioral tasks, where specific neural patterns reveal dynamic evolutions corresponding to distinct behaviors of interest. However, empirically addressing this scientific question is challenging due to neural population activities in various brain regions exhibiting mixed selectivity (Sani et al., 2021; Hasnain et al., 2023), responding robustly to multiple behaviors of interest. We further verify this finding through an empirical study on the dorsal cortex of head-fixed mice (Musall et al., 2019) (shown in Figure 1).
Figure 1: Empirical study on dorsal cortex neural recordings of a mouse in a visual decision-making task. (A) The Brain Atlas map (Lein et al., 2007). (B) Neural signals in various brain regions (SSp, MOs, and VIS) exhibit mixed selectivity in behavior of interest decoding. “Levers”, “Spouts”, “Paw-(x)”, and “Jaw” are four behaviors of interest. $\mathbf{cvR}^{2}$ is short for cross-validation coefficient of determination. The higher, the better.
To tackle this issue, we propose a method - Exploring Behavior-Relevant and Interpretable Neural Dynamics with Generative Diffusion Models - (“BeNeDiff”). We first employ a neural latent variable model (LVM) to identify orthogonal and disentangled neural latent subspace. This is achieved through a semi-supervised variational autoencoder, which integrates behavioral labels to rotate the subspace. Subsequently, our main idea is to explore the neural dynamics of each latent factor in the learnt subspace for distinct quantifications of the behaviors of interest. However, such a workflow is non-trivial since naïve latent manipulation produces samples not conform to the original distribution, leading to mapped video-based behavioral data that loses its validity (we further detail this part in Method Section 3.2.1).
@@
We conduct experiments to verify the efficacy of BeNeDiff on a widefield calcium imaging dataset, where a head-fixed mouse performs a visual decision-making task across multiple sessions (Musall et al., 2018, 2019). The neural subspace in BeNeDiff exhibits high levels of disentanglement and neural reconstruction quality, as evidenced by multiple quantitative metrics. By guiding the diffusion model to activate individual latent factors, we verify that the neural dynamics within the disentangled subspace provide interpretable and selective quantifications of the behaviors of interest (e.g., paw movements). These results advance our understanding of neuro-behavioral relationships through the identification of fine-grained behavioral subspaces and the uncovering of disentangled neural dynamics.
To highlight our major contributions: (1) This is the first work to explore wide-field imaging of the dorsal cortex of mice during a decision-making task using neural subspace analysis, rather than merely performing neuro-behavior decoding. We uncover disentangled neural representations for various behaviors. (2) To visualize the behavior dynamics within a disentangled neural subspace, we develop a novel VDM-based interpretation tool that faithfully reflects behavior-related neural dynamics. It is essential to interpret the meaning of each neural latent dimension as well as the behavior dynamics it encodes.
@@ 2.1 Problem Formulation
We first provide the notations of the paired neuro-behavioral observations. The single-trial neural population activities are denoted as $\mathbf{X}=\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}\right]^{\top}\in\mathbb{R}^{L\times N}$, where $L$ is the trial length (i.e., number of time bins), $N$ is the number of observed neural signals. The behavioral video frames are denoted as $\mathbf{Y}=\left[\mathbf{Y}_{1},\ldots,\mathbf{Y}_{L}\right]^{\top}\in\mathbb{R}^{L\times H\times W}$, where $H$, $W$ are the height and width of the compressed behavior video frames. We extract behavior labels $\mathbf{U}=\left[\mathbf{u}_{1},\ldots,\mathbf{u}_{L}\right]^{\top}\in\mathbb{R}^{L\times B}$ from the video frames using a behavior LVM (Whiteway et al., 2021). $B$ is the number of the behavior of interest.
@@ 3 Methods
Figure 2: Schematic diagram of neural dynamics interpretation with BeNeDiff. We first employ a neural LVM to identify a disentangled neural latent subspace (the left part). Then, we leverage video diffusion models (VDMs) to generate behavior videos predicted to activate the variance of individual latent factors along the single-trial trajectory, providing interpretable quantifications of neural dynamics with respect to the behaviors of interest.
@@ 3.1 Behavior-Relevant and Disentangled Neural Latent Subspace Learning
Drawing inspiration from recent progress in the field of neural LVMs (Kingma et al., 2014; Klys et al., 2018), we propose to employ a VAE to learn a disentangled neural subspace. The neural data $\mathbf{X}$ usually contains a good amount of information other than behavior (Hasnain et al., 2023), thus an unsupervised disentangled VAE won’t effectively discover disentangled subspace with behavior only. Therefore, we introduce behavior labels $\mathbf{U}$ to inform the VAE to learn a latent subspace that better accounts for the variance related to behavior. We note that this technique is widely adopted in previous neuro-behavioral analysis works (Wang et al., 2024; Schneider et al., 2023; Gondur et al., 2023). Notably, to enforce the disentanglement in the latent subspace, we propose to incorporate a total-correlation (TC) penalty term (Chen et al., 2018) to enforce the VAE to find statistically independent latent factors in the semi-supervised setting. The VAE optimizes the following evidence lower bound (ELBO) (MacKay, 2003):
@@ 3.2.1 Downside of Latent Manipulation for Interpreting Neural Dynamics
As for testifying the neural dynamics of a single disentangled latent factor $\mathbf{z}^{(d)}$ on the behavioral videos $\mathbf{Y}$, a straightforward attempt is to train a neural-net model to approximate the posterior distribution $p(\mathbf{Y}\mid\mathbf{Z})$ and then perform latent manipulation on each single latent factor. There are two major techniques to perform latent manipulation. The first is a naïve manipulation. Formally, the single latent factor manipulation is defined as $\tilde{\mathbf{z}}^{(d)}=\mathbf{z}^{(d)}+\boldsymbol{\eta},\;\boldsymbol{\eta}\sim\mathcal{U}(-a,a)$, where $a$ is the half-range of the uniform distribution. The second method uses classifier-free guidance (Ho and Salimans, 2022), where we allow the activated latent factor $\mathbf{z}^{(d)}$ to evolve while setting the values of the other latent factors to those from the first frame of the trial. We denote the overall neural trajectory after manipulating the latent factor $\tilde{\mathbf{z}}^{(d)}$ as $\tilde{\mathbf{Z}}$. Through either of these manipulation methods, $\tilde{\mathbf{Z}}$ may not necessarily conform to the overall neural latent trajectory distribution $p(\mathbf{Z})$, thus making the output behavioral videos hard for visualization or interpretation.
@@ 3.2.2 Behavioral Video Generation for Neural Dynamics Interpretation
So here we employ the video diffusion models (VDMs) to explore factor-wise neural dynamics through a generative manner, which is capable of maintaining temporal consistency and behavioral dynamics across frames. The primary goal is to perform behavior data generation conditioned on activating a single latent factor along the neural trajectory. Thus the resulting behavior video can provide interpretable quantifications of the neural dynamics of factor $\mathbf{z}^{(d)}$. Technically, we implement classifier-guidance (Kawar et al., 2022). By Bayes rule, we obtain the following posterior density and gradient (Mardani et al., 2023):
@@
For the calculation of the guidance term, we first train a linear encoder as the classifier from the behavior video data to the neural latent variables of the learned semi-supervised VAE subspace. We denote the estimated neural latent trajectories as $\hat{\mathbf{Z}}_{t}=\left[\hat{\mathbf{z}}_{t,1},\ldots,\hat{\mathbf{z}}_{t,L}\right]^{\top}\in\mathbb{R}^{L\times D}$, in which:
@@
The class labels $\mathbf{c}\in\left\{\mathbf{c}^{(1)},\mathbf{c}^{(2)},\ldots,\mathbf{c}^{(D)}\right\}$, in which $\mathbf{c}^{(d)}$ is a one-hot column vector with a one at the $d$-th dimension and zeros elsewhere. Drawing inspiration from Noise-Contrastive Estimation (Gutmann and Hyvärinen, 2010), our guidance objective of the activation of latent factor $d$-th is formulated as maximizing the variance of the trajectory $\hat{\mathbf{z}}^{(d)}_{t}$ while minimizing the variance of the other latent factor trajectories in $\hat{\mathbf{Z}}_{t}$:
$\log p_{\boldsymbol{\lambda}}\left(\mathbf{c}^{(d)}\middle|\mathbf{Y}_{t}\right)=\log\left[\frac{\exp\left(f^{+}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}}_{t},\mathbf{c}^{(d)}\right)/\tau\right)}{\exp\left(f^{+}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}}_{t},\mathbf{c}^{(d)}\right)/\tau\right)+\sum_{k=1}^{K}\exp\left(f^{-}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}}_{t},\mathbf{c}^{(d)}\right)/\tau\right)}\right],$ (8)
where $f^{+}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}}_{t},\mathbf{c}^{(d)}\right)=\operatorname{Var}\left(\hat{\mathbf{Z}}_{t}\right)\mathbf{c}^{(d)}$ calculates the variance of the selected latent factor and $f^{-}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}}_{t},\mathbf{c}^{(d)}\right)=\operatorname{Var}\left(\hat{\mathbf{Z}}_{t}\right)\mathbf{c}^{(j)},\;j\sim\operatorname{Uniform}(\{1,2,\ldots,D\}\backslash\{d\})$ calculates the variance of another sampled latent factor’s trajectory. $\operatorname{Var}\left(\hat{\mathbf{Z}}_{t}\right)\in\mathbb{R}^{1\times D}$ is a row vector where each element is the variance of every latent factor along the neural trajectory. $\tau$ is the temperature parameter. $K$ is a hyperparameter controlling the number of sampled negative samples at each iteration.
@@ 4 Related Works
Disentangled Latent Subspace Learning.  Neural LVMs is a fundamental framework which posits that single-trial neural population activities rely on low-dimensional “neural manifolds” (Gallego et al., 2018; Mitchell-Heggs et al., 2023; Li et al., 2023a) and their extracted latent variables are successful in describing single-trial neural activities (Li et al., 2024c, 2022, d, a). Learning disentangled latent variables that uncover statistically independent latent factors (Chen et al., 2018; Kim and Mnih, 2018) can provide enhanced robustness, interpretability, and controllability. Typically, this type of work involves adding auxiliary regularizer terms to enhance orthogonality (Mathieu et al., 2019) and reduce the total correlation (Chen et al., 2018) among the latent factors. In the field of neuroscience, there have been studies focusing on the disentanglement of latent subspace within rich behavioral data (Whiteway et al., 2021; Shi et al., 2021). However, our work is the first to discover interpretable and disentangled latent subspaces of wide-field imaging data.
Generative Diffusion Models.   In recent years, diffusion models have achieved great success in generating high-quality images due to their expressivity and flexibility (Ho et al., 2020; Song et al., 2020a, b; Vahdat et al., 2021). Moreover, for the more challenging task of video generation, there have been several explorations using diffusion models to address it. From a modeling perspective, the key concern is how to maintain temporal dynamics (Li et al., 2021) and consistency across frames. Most existing works (Ho et al., 2022b, a) extend the 2D U-Net architecture (Ronneberger et al., 2015) to a 3D framework by considering the time axis. In this 3D framework, convolutions are performed in both spatial and temporal dimensions. Additionally, recent studies in neural computation have leveraged generative diffusion models to tackle domain-specific tasks, such as neural distribution alignment (Wang et al., 2024) and decoding visual stimulus from brain activities (Sun et al., 2024b, a). Our work is the first to employ generative diffusion models for analyzing neuro-behavioral data relationships.

### v3

@@ Abstract
Understanding the neural basis of behavior is a fundamental goal in neuroscience. Current research in large-scale neuro-behavioral data analysis often relies on decoding models, which quantify behavioral information in neural data but lack details on behavior encoding. This raises an intriguing scientific question: “how can we enable in-depth exploration of neural representations in behavioral tasks, revealing interpretable neural dynamics associated with behaviors”. However, addressing this issue is challenging due to the varied behavioral encoding across different brain regions and mixed selectivity at the population level. To tackle this limitation, our approach, named “BeNeDiff”, first identifies a fine-grained and disentangled neural subspace using a behavior-informed latent variable model. It then employs state-of-the-art generative diffusion models to synthesize behavior videos that interpret the neural dynamics of each latent factor. We validate the method on multi-session datasets containing wide-field calcium imaging recordings across multiple brain regions of the dorsal cortex. By guiding the diffusion model to activate individual latent factors, we verify that the neural dynamics of latent factors in the disentangled neural subspace provide interpretable quantifications of the behaviors of interest across multiple brain regions. Meanwhile, the neural subspace in BeNeDiff demonstrates high disentanglement and neural reconstruction quality. Our codes are available at https://github.com/BRAINML-GT/BeNeDiff.
@@ 1 Introduction
Understanding and elucidating the complex interrelationships between behavioral data and neural population activity is a long-standing goal in systems neuroscience (Batty et al., 2019; Gomez-Marin et al., 2014; Krakauer et al., 2017; Berman, 2018). Exploring the neural basis of behavior not only deepens our basic knowledge of brain functions but also establishes a foundation for developing improved treatments for psychiatric and neurological conditions (Vieira et al., 2017; Ibáñez et al., 2018). Significant progress has been achieved in developing computational toolkits for neuro-behavioral decoding by using behavior video data (Whiteway et al., 2021; Batty et al., 2019; Musall et al., 2019; Stringer et al., 2019). These methods perform region-based behavior decoding to map neural activity across multiple brain regions of the dorsal cortex to the behaviors from the videos. However, these methods only quantify how much behavioral information is encoded in neural populations, but do not reveal the details of such encoding. There has been markedly less focus, with cortex-wide signals, on enabling in-depth exploration of neural activities during behavioral tasks, where specific neural patterns reveal dynamic evolutions corresponding to distinct behaviors of interest. However, empirically addressing this scientific question is challenging due to neural population activities in various brain regions exhibiting mixed selectivity (Sani et al., 2021; Hasnain et al., 2023), responding robustly to multiple behaviors of interest. We further verify this finding through an empirical study across three brain regions on the dorsal cortex of head-fixed mice (Musall et al., 2019) (shown in Figure 1).
Figure 1: Empirical study across multiple brain regions of dorsal cortex neural recordings of a mouse in a visual decision-making task. (A) The Brain Atlas map (Lein et al., 2007). (B) Neural signals in various brain regions (SSp, MOs, and VIS) exhibit mixed selectivity in behavior of interest decoding. “Levers”, “Spouts”, “Paw-(x)”, and “Jaw” are four behaviors of interest. $\mathbf{cvR}^{2}$ is short for cross-validation coefficient of determination. The higher, the better.
To tackle this issue, we propose a method - Exploring Behavior-Relevant and Interpretable Neural Dynamics with Generative Diffusion Models - (“BeNeDiff”). We first employ a neural latent variable model (LVM) to identify orthogonal and disentangled neural latent subspace. This is achieved through a semi-supervised variational autoencoder, which integrates behavioral labels to rotate the subspace. Subsequently, our main idea is to explore the neural dynamics of each latent factor in the learned subspace for distinct quantifications of the behaviors of interest. However, such a workflow is non-trivial since naïve latent manipulation produces samples not conform to the original distribution, leading to mapped video-based behavioral data that loses its validity (we further detail this part in Method Section 3.2.1).
@@
We conduct experiments to verify the efficacy of BeNeDiff on a widefield calcium imaging dataset, where a head-fixed mouse performs a visual decision-making task across multiple sessions (Musall et al., 2018, 2019). The neural subspace in BeNeDiff exhibits high levels of disentanglement and neural reconstruction quality, as evidenced by multiple quantitative metrics. By guiding the diffusion model to activate individual latent factors, we verify that the neural dynamics within the disentangled subspace provide interpretable and selective quantifications of the behaviors of interest (e.g., paw movements) across multiple brain regions. These results advance our understanding of neuro-behavioral relationships through the identification of fine-grained behavioral subspaces and the uncovering of disentangled neural dynamics.
To highlight our major contributions: (1) This is the first work to explore wide-field imaging across multiple brain regions of the dorsal cortex of head-fixed mice during a decision-making task using neural subspace analysis, rather than merely performing neuro-behavior decoding. We uncover disentangled neural representations for various behaviors. (2) To visualize the behavior dynamics within a disentangled neural subspace of each brain region, we develop a novel VDM-based interpretation tool that faithfully reflects behavior-related neural dynamics. It is essential to interpret the meaning of each neural latent dimension as well as the behavior dynamics it encodes.
@@ 2.1 Problem Formulation
We first provide the notations of the paired neuro-behavioral observations. The single-trial neural population activities are denoted as $\mathbf{X}=\left[\mathbf{x}_{1},\ldots,\mathbf{x}_{L}\right]^{\top}\in\mathbb{R}^{L\times N}$, where $L$ is the trial length (i.e., number of time bins), $N$ is the number of observed neural signals. The behavioral video frames are denoted as $\mathbf{Y}=\left[\mathbf{Y}_{1},\ldots,\mathbf{Y}_{L}\right]^{\top}\in\mathbb{R}^{L\times H\times W}$, where $H$, $W$ are the height and width of the compressed behavior video frames. We extract behavior labels $\mathbf{U}=\left[\mathbf{u}_{1},\ldots,\mathbf{u}_{L}\right]^{\top}\in\mathbb{R}^{L\times B}$ from the video frames using a behavior LVM (Whiteway et al., 2021). $B$ is the number of the behavior.
@@ 3 Methods
Figure 2: Schematic diagram of neural dynamics interpretation with BeNeDiff. We first employ a neural LVM to identify a disentangled neural latent subspace (the left part). Then, we train a linear neural encoder to map behavior video frames to neural trajectories. We use video diffusion models (VDMs) to generate behavior videos guided by the neural encoder, based on the objective of activating the variance of individual latent factors along the single-trial trajectory. This approach provides interpretable quantifications of neural dynamics in relation to the behaviors of interest.
Then, we train a linear neural encoder from the behavior video frames to the neural trajectories. We leverage video diffusion models (VDMs) to generate behavior videos guided by the neural encoder, based on the objective of activating the variance of individual latent factors along the single-trial trajectory, providing interpretable quantifications of neural dynamics with respect to the behaviors of interest.
@@ 3.1 Behavior-Relevant and Disentangled Neural Latent Subspace Learning
Drawing inspiration from recent progress in the field of neural LVMs (Kingma et al., 2014; Klys et al., 2018), we employ a VAE to learn a disentangled neural subspace. The neural data $\mathbf{X}$ usually contains a good amount of information other than behavior (Hasnain et al., 2023), thus an unsupervised disentangled VAE won’t effectively discover disentangled subspace with behavior only. Therefore, we introduce behavior labels $\mathbf{U}$ to inform the VAE to learn a latent subspace that better accounts for the variance related to behavior. We note that this technique is widely adopted in previous neuro-behavioral analysis works (Wang et al., 2024; Schneider et al., 2023; Gondur et al., 2023). Notably, to enforce the disentanglement in the latent subspace, we incorporate a total-correlation (TC) penalty term (Chen et al., 2018) to enforce the VAE to find statistically independent latent factors in the semi-supervised setting. The VAE optimizes the following evidence lower bound (ELBO) (MacKay, 2003):
@@ 3.2.1 Downside of Latent Manipulation for Interpreting Neural Dynamics
As for testifying the neural dynamics of a single disentangled latent factor $\mathbf{z}^{(d)}$ on the behavioral videos $\mathbf{Y}$, a straightforward attempt is to train a neural-net model to approximate the posterior distribution $p(\mathbf{Y}\mid\mathbf{Z})$ and then perform latent manipulation on each single latent factor. There are two major techniques to perform latent manipulation. The first is a naïve manipulation. This method manipulates a single subspace $\mathbf{z}^{(d)}$ while keeping the non-target latent factors fixed at arbitrary values. It then observes how the manipulation affects $\mathbf{Y}$. The induced changes in the videos reveal the dynamics encoded by $\mathbf{z}^{(d)}$. The second method uses classifier-free guidance (Ho and Salimans, 2022), where we allow the activated latent factor $\mathbf{z}^{(d)}$ to evolve while fixing non-target latent factors to arbitrary values. However, setting arbitrary values without knowing the true distributions of non-target subspaces can lead to unnatural distortions in generated videos, complicating the visualization and interpretation of genuine animal behavioral dynamics.
@@ 3.2.2 Behavioral Video Generation for Neural Dynamics Interpretation
So here we employ the video diffusion models (VDMs) to explore factor-wise neural dynamics through a generative manner, which is capable of maintaining temporal consistency and behavioral dynamics across frames. The primary goal is to perform behavior data generation conditioned on activating a single latent factor along the neural trajectory. Thus the resulting behavior video can provide interpretable quantifications of the neural dynamics of factor $\mathbf{z}^{(d)}$. Specifically, we implement classifier guidance (Kawar et al., 2022). By Bayes rule, we obtain the following posterior density and gradient (Mardani et al., 2023):
@@
For the calculation of the guidance term, we first train a linear neural encoder as the classifier from the behavior video data to the neural latent variables of the learned semi-supervised VAE subspace. We denote the estimated neural latent trajectories as $\hat{\mathbf{Z}}_{t}=\left[\hat{\mathbf{z}}_{t,1},\ldots,\hat{\mathbf{z}}_{t,L}\right]^{\top}\in\mathbb{R}^{L\times D}$, in which:
@@
The class labels $\mathbf{c}\in\left\{\mathbf{c}^{(1)},\mathbf{c}^{(2)},\ldots,\mathbf{c}^{(D)}\right\}$, in which $\mathbf{c}^{(d)}$ is a one-hot column vector with a one at the $d$-th dimension and zeros elsewhere. Drawing inspiration from Noise-Contrastive Estimation (Gutmann and Hyvärinen, 2010), our guidance objective of the activation of latent factor $d$-th is formulated as maximizing the variance of the trajectory $\hat{\mathbf{z}}^{(d)}$ while minimizing the variance of the other latent factor trajectories in $\hat{\mathbf{Z}}$:
$\log p_{\boldsymbol{\lambda}}\left(\mathbf{c}^{(d)}\middle|\mathbf{Y}_{t}\right)=\log\left[\frac{\exp\left(f^{+}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}},\mathbf{c}^{(d)}\right)/\tau\right)}{\exp\left(f^{+}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}},\mathbf{c}^{(d)}\right)/\tau\right)+\sum_{k=1}^{K}\exp\left(f^{-}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}},\mathbf{c}^{(d)}\right)/\tau\right)}\right],$ (8)
where $f^{+}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}},\mathbf{c}^{(d)}\right)=\operatorname{Var}\left(\hat{\mathbf{Z}}\right)\mathbf{c}^{(d)}$ calculates the variance of the selected latent factor and $f^{-}_{\boldsymbol{\lambda}}\left(\hat{\mathbf{Z}},\mathbf{c}^{(d)}\right)=\operatorname{Var}\left(\hat{\mathbf{Z}}\right)\mathbf{c}^{(j)},\;j\sim\operatorname{Uniform}(\{1,2,\ldots,D\}\backslash\{d\})$ calculates the variance of another sampled latent factor’s trajectory. $\operatorname{Var}\left(\hat{\mathbf{Z}}\right)\in\mathbb{R}^{1\times D}$ is a row vector where each element is the variance of every latent factor along the neural trajectory. $\tau$ is the temperature parameter. $K$ is a hyperparameter controlling the number of sampled negative samples at each iteration.
@@ 4 Related Works
Disentangled Latent Subspace Learning.  Neural LVMs is a fundamental framework which posits that single-trial neural population activities rely on low-dimensional “neural manifolds” (Gallego et al., 2018; Mitchell-Heggs et al., 2023; Li et al., 2023a) and their extracted latent variables are successful in describing single-trial neural activities (Li et al., 2024c, 2022, d; Liu et al., 2021, 2022; Li et al., 2024a). Learning disentangled latent variables that uncover statistically independent latent factors (Chen et al., 2018) can provide enhanced robustness, interpretability, and controllability. Typically, this type of work involves adding auxiliary regularizer terms to enhance orthogonality (Mathieu et al., 2019) and reduce the total correlation (Chen et al., 2018) among the latent factors. In the field of neuroscience, there have been studies focusing on the disentanglement of latent subspace within rich behavioral data (Whiteway et al., 2021; Shi et al., 2021). However, our work is the first to discover interpretable and disentangled latent subspaces of wide-field imaging data.
Generative Diffusion Models.   In recent years, diffusion models have achieved great success in generating high-quality images due to their expressivity and flexibility (Ho et al., 2020; Song et al., 2020a, b; Vahdat et al., 2021). Moreover, for the more challenging task of video generation, there have been several explorations using diffusion models to address it. From a modeling perspective, the key concern is how to maintain temporal dynamics and consistency across frames. Most existing works (Ho et al., 2022b, a) extend the 2D U-Net architecture (Ronneberger et al., 2015; Song et al., 2024) to a 3D framework by considering the time axis. In this 3D framework, convolutions are performed in both spatial and temporal dimensions. Additionally, recent studies in neural computation have leveraged generative diffusion models to tackle domain-specific tasks, such as neural distribution alignment (Wang et al., 2024) and decoding visual stimulus from brain activities (Sun et al., 2024b, a). Our work is the first to employ generative diffusion models for analyzing neuro-behavioral data relationships.